OpenStudy (anonymous):

A sample of ideal gas is in a sealed container. The pressure of the gas is 505torr, and the temperature is 43 degree celsius. If the temperature changes to 93 degree celsius with no change in volume or amount of gas, what is the new pressure,P2 , of the gas inside the container?

OpenStudy (rogue):

First, convert your temperatures from Celsius to kelvin. k = C + 273.15. Next, use that formula (I have no idea what its name is)\[\frac {P_1}{T_1} = \frac {P_2}{T_2}\]

OpenStudy (anonymous):

I dont know how to use that formula

OpenStudy (anonymous):

-___-

OpenStudy (rogue):

-___-

OpenStudy (rogue):

Okays. P1 and T1 are your initial pressures and temperatures. P2 and T2 are your final pressures and temperatures. P1 = 505 torr T1 = 43 C = 316 k P2 = ? T2 = 93 C = 366 k Plug them into your formula\[\frac {505 \space torr}{316 \space k} = \frac {P_2}{366 \space k}\]\[P_2 = \frac {505 \space torr \times 366 \space k}{316 \space k} \approx 584.905 \space torr\]

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

can you leave me with the second part please

OpenStudy (anonymous):

Using the same sample of gas (P1 = 505 torr ,T1 = 43 degree C.), we wish to change the pressure to 5050 torr with no accompanying change in volume or amount of gas. What temperature T2 , in Celsius, is needed to reach this pressure?

OpenStudy (rogue):

Alright, instead of solving this one via the formula, I explain the concept behind it. Pressure and temperature are directly proportional. When volume and moles are kept constant, if pressure changes by a factor x, the temperature will change by the same factor x. Let's say the initial pressure & temp of a sample of gas is 50 torr and 400 k. If the pressure is doubled to 100 torr, the temp will also double along with the pressure to 800 k. In your case, the pressure increases by 10 times, from 505 torr to 5050 tor. So the final temperature will also increase by 10 times. So the final temperature will be 316 k x 10 = 3160 k. If you want that back in Celsius, subtract 273 and you get the final temp is 2887 C.